# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v1

@@ 5.1 Directed Graph Generation Performance
Table 1 reports the results on two synthetic datasets: ER-DAG and SBM. In both settings, Directo consistently achieves the best trade-off between sample quality and validity. For ER-DAG, it achieves the highest validity score using directed RRWP as positional encoding, with a V.U.N. ratio of 94%. Notably, while LayerDAG enforces acyclicity by design, it fails to capture the ER structure of the target graph distribution, as evidenced by its low V.U.N. score (21.5%; see Tab. 13 for details). In the SBM case, Directo also successfully captures the distribution and consistently outperforms the baselines, achieving up to 95.5% in V.U.N. ratio and consistently low average graph statistics ratios. In Appendix H, we also provide results for two additional synthetic datasets as well as a study on the effect of dataset size, which shows that larger datasets lead to improved performance but also incur higher computational cost and longer training times.
@@
Table 2 shows the performance for the two real-world datasets. In TPU tiles, all Directo variants significantly outperform DAG generation baselines in terms of structure-aware metrics. The ratio scores for Directo are consistently and substantially lower than those observed in MLE, and especially LayerDAG. This indicates a more realistic distribution. Again, Directo variants achieve higher V.U.N. than all baselines, with the exception of LayerDAG, which benefits from enforcing acyclicity, the only validity constraint evaluated in this case. For Visual Genome, results reinforce the versatility of Directo. Directo RRWP achieves the highest V.U.N. scores, indicating a superior ability to generate diverse and novel graphs while maintaining structural fidelity.
5.2 Dual Attention Ablation
@@
We now analyze the influence of the dual attention mechanism on the performance of Directo. In Fig. 3, we compare the V.U.N metric and the average graph statistics ratios ("Ratio") score across both synthetic datasets (ER-DAG and SBM), highlighting the impact of dual attention on generation quality and graph realism. We observe that the dual attention mechanism consistently improves generative performance, regardless of the positional encoding it is combined with. Notably, even in the absence of any positional encoding ("No PE"), dual attention still achieves non-zero V.U.N., highlighting its capacity to independently capture directionality-relevant information. See Appendix H.1 for the full ablation tables, both for Directo and Directo-DD.
TPU Tiles
Training set 0.0003 0.0003 0.0007 0.0006 0.0002 1.0 100 100 0 0
D-VAE OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
DeFoG 0.0099 $\pm$ 0.0012 0.0794 $\pm$ 0.0023 0.0042 $\pm$ 0.0027 0.0040 $\pm$ 0.0005 0.0017 $\pm$ 0.0003 63.7 $\pm$ 2.6 86.3 $\pm$ 2.2 87.7(∗) $\pm$ 5.4 93.3 $\pm$ 2.5 72 $\pm$ 2.4
Directo-DD MagLap 0.0115 $\pm$ 0.0035 0.0703 $\pm$ 0.0033 0.0103 $\pm$ 0.0021 0.0076 $\pm$ 0.0007 0.0043 $\pm$ 0.0014 64.3 $\pm$ 5.3 86.5 $\pm$ 5.1 90.5 $\pm$ 3.3 100 $\pm$ 0.0 77 $\pm$ 7.0
Directo RRWP 0.0133 $\pm$ 0.0025 0.0859 $\pm$ 0.0072 0.0136 $\pm$ 0.0075 0.0086 $\pm$ 0.0010 0.0038 $\pm$ 0.0009 75.4 $\pm$ 8.1 97 $\pm$ 1.0 81.8(∗) $\pm$ 4.5 96.5 $\pm$ 1.2 77 $\pm$ 2.9
Visual Genome
Training set 0.0018 0.0030 0.0000 0.0072 0.0036 1.0 100 100 0 0
MLE 0.0607 $\pm$ 0.0036 0.0474 $\pm$ 0.0023 0.5342 $\pm$ 0.0830 0.0535 $\pm$ 0.0015 0.0399 $\pm$ 0.0017 17.0 $\pm$ 0.6 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
Directo-DD RRWP 0.0424 $\pm$ 0.0043 0.0413 $\pm$ 0.0037 $0$$\pm$ 0.0000 0.0132 $\pm$ 0.0053 0.0074 $\pm$ 0.0008 15.3 $\pm$ 0.8 72.7 $\pm$ 3.9 100 $\pm$ 0.0 100 $\pm$ 0.0 72.7 $\pm$ 3.9
Directo-DD MagLap 0.0245 $\pm$ 0.0022 0.0359 $\pm$ 0.0036 $0$$\pm$ 0.0000 0.0163 $\pm$ 0.0026 0.0086 $\pm$ 0.0011 7.6 $\pm$ 0.7 86.5 $\pm$ 5.1 100 $\pm$ 0.0 100 $\pm$ 0.0 61.9 $\pm$ 4.4
Directo RRWP 0.0494 $\pm$ 0.0021 0.0425 $\pm$ 0.0029 $0$$\pm$ 0.0000 0.0226 $\pm$ 0.0075 0.0228 $\pm$ 0.0058 12.8 $\pm$ 0.6 86 $\pm$ 4.5 98.4 $\pm$ 1.7 99.3 $\pm$ 0.7 83.8 $\pm$ 4.3
Directo MagLap 0.0180 $\pm$ 0.0024 0.0302 $\pm$ 0.0040 $0$$\pm$ 0.0000 0.0142 $\pm$ 0.0021 0.0099 $\pm$ 0.0013 6.2 $\pm$ 0.5 67.6 $\pm$ 3.6 99.7 $\pm$ 0.6 99.7 $\pm$ 0.6 67.0 $\pm$ 4.3
@@
5.3 Positional Encodings Ablation
We also evaluate the sensitivity of Directo to different positional encodings, including a direction-agnostic baseline ("Lap"). In Fig. 3, we observe consistent trends across datasets: integrating positional encodings improves V.U.N. and Ratio, mirroring the undirected setting; moreover, direction-aware encodings outperform agnostic ones, supporting our modeling hypothesis. Among the encodings tested, Directed RRWP achieves the best overall performance on ER-DAG, and MultiMagLapPE with $Q=10$ on SBM. Nevertheless, we remark that RRWP demonstrates clear superiority in both scalability and runtime efficiency (see Sec. G.3). Full tables for both Directo and Directo-DD can be found in Appendix H.2.
@@ 6 Related work
Generative models that were initially adapted to the graph setting include auto-regressive methods you2018GraphRNN ; liao2019gran , which grow the graph progressively by inserting nodes and edges, but require node ordering for feasibility. On the other hand, one-shot graph generative models such as VAEs kipf2016variational ; simonovsky2018graphvae ; jin2019junctiontreevae ; ma2021gfvae ; vignac2022topn , GANs de2018molgan ; krawczuk2021gggan ; martinkus2022SPECTRE , and normalizing flows kaushalya2019graphnvp ; wehenkel2021graphnormalizingflows ; lippe2021categorical ; luo21graphdf can directly generate full graphs in a single forward pass, thereby eliminating the need for predefined node ordering. Graph diffusion models also belong to this class, with the initial approaches consisting of adaptations of continuous state-space discrete-time diffusion frameworks sohl2015deep ; ho2020denoising ; song2020denoising to the graph settin. These include works such as EDP-GNN niu2020permutation , DGSM luo2021predicting , and GeoDiff xu2022geodiff . Later, the discrete state-space diffusion framework austin2021structured was adopted, with DiGress vignac2023digress and MCD haefeli2022diffusion operating in discrete time. On the other hand, methods that operate in continuous time campbell2022continuous were introduced, including GDSS jo2022score and GruM jo2024GruM in continuous state-spaces and DisCo xu2024disco and Cometh siraudin2024cometh for discrete state-spaces. Recent approaches combine the sequential modeling of auto-regressive methods with the global denoising of diffusion zhao2024pard . Finally, DeFoG qin2025defog recently achieves state-of-the-art performance by leveraging discrete flow matching campbell2024generative ; gat2024discrete .
@@
The vast majority of works in this area focus solely on directed acyclic graph (DAG) generation, with research including autoencoders zhang2019dvae , autoregressive methods li2022graphpnas and, more recently, combining diffusion in different steps of an auto-regressive process li2024layerdag . However, they require the input DAGs to be topologically ordered. Other approaches asthan2024NAS ; an2023diffusionnag use ideas from discrete diffusion to perform Neural Architecture Search (NAS), but they are tailored to the specific application and, therefore, to DAGs. Finally, the work in law2024directedheatkernels is the first to propose a general directed graph generation method, approaching it from an auto-regressive formulation. Unfortunately, to the best of our knowledge, its implementation is not publicly available, which limits its accessibility for study and evaluation.
@@ 7 Conclusion and future directions
We propose Directo, a novel discrete flow matching-based method for effective directed graph generation. To achieve this, we combine directionality-aware positional encodings with a dual-attention mechanism that explicitly handles source-to-target and target-to-source dependencies. We additionally establish a new benchmarking framework tailored to digraphs, addressing the current lack of standardized evaluation in the field. Empirical results demonstrate that our model significantly outperforms existing baselines in a range of synthetic and real-world datasets. Our approach successfully learns to generate directed structures and preserve critical properties, such as acyclicity, without the need for explicit constraints. Looking forward, scalability, a deeper understanding of the role of permutation equivariance, and further guidance of the generative process remain promising directions of research.
@@ Appendix A Dual Attention
To incorporate edge features into the attention mechanism, we apply FiLM-style modulation using additive and multiplicative projections $\bm{W}_{\text{mul}},\bm{W}_{\text{add}}$:
@@
where $X_{\to y}$ and $E_{\to y}$ are PNA layers corso2020principal that given node features $\bm{X}\in\mathbb{R}^{B\times N\times d_{x}}$ or edge features $\bm{E}\in\mathbb{R}^{B\times N\times N\times d_{e}}$ and a learnable weight matrix $\bm{W}_{\text{PNA}}$, computes
@@ Appendix B Positional Encodings
To retain directional information while preserving desirable spectral properties, authors in geisler2023transformers propose the Magnetic Laplacian, which introduces a complex-valued phase encoding into the adjacency matrix. Specifically, the (unnormalized and normalized) Magnetic Laplacians are given by:
@@
Recent work by huang2025good extends the Magnetic Laplacian by introducing a multi-$q$ formulation, which considers as positional encoding stacking together the eigenvalues and eigenvectors of $Q$ distinct Magnetic Laplacians with potentials $q_{1},\ldots,q_{Q}$. This approach enables the recovery of the bidirectional walk profile, a generalization of walk counting in undirected graphs that captures a broader range of directional relationships.
@@
For directed graphs, it is possible to consider $\bm{T}=\bm{A}\bm{D}^{-1}_{out}$, where we take into account the degree matrix of the outgoing edges $\bm{D}^{-1}_{out}$ geisler2023transformers . The $K$-step transition probabilities are then captured by powers of $\bm{T}$, producing the sequence $[\bm{I},\bm{T},\bm{T}^{2},\ldots,\bm{T}^{K-1}]$.
Additionally, in the directed setting, authors in geisler2023transformers suggest it is also informative to model reverse random walks, starting from incoming edges. For this, we can define the reverse transition matrix $\bm{R}=\bm{A}^{\top}\bm{D}_{\text{in}}^{-1}$, where $\bm{D}_{\text{in}}$ is the in-degree matrix. This would represent the likelihood of arriving at a given edge starting from another through $k$-step random walks. Analogously, we compute $[\bm{I},\bm{R},\bm{R}^{2},\ldots,\bm{R}^{K-1}]$.
To have our final positional encoding, following geisler2023transformers , we ensure that graphs are not nilpotent and that the probabilities encoded in the matrices sum up to 1 by adding self-loops to sink nodes during pre-processing. Finally, the full positional encoding concatenates both forward and reverse walk features for a pre-defined $K$, yielding:
@@ Appendix C Sampling Optimization in Directo
The decoupling of training and sampling in discrete flow matching enables principled, training-free modifications to the sampling procedure, offering a flexible design space to enhance generation quality across digraph distributions qin2025defog . In this section, we detail the three core parameters that govern sampling behavior: (i) time distortion functions, (ii) target guidance intensity, and (iii) stochasticity strength. Each plays a distinct role in shaping the denoising trajectory and contributes to different aspects of generation quality and structural fidelity.
@@
We consider the same five representative time distortion functions from qin2025defog that yield diverse distributions for $t^{\prime}$:
@@
The final sampling-time parameter is the stochasticity coefficient $\eta$. This parameter scales an auxiliary rate matrix $R_{t}^{\text{DB}}$ that satisfies the detailed balance condition campbell2024generative and adds it to the optimal rate matrix $R_{t}^{*}$:
@@
Together, these sampling parameters constitute a flexible, principled toolkit for tailoring the sampling procedure to diverse graph domains and structural requirements. In our case, we optimize these parameters at sampling time to improve the performance of Directo. In particular, we individually optimize each of them, and perform the final sampling using the best performing parameters in terms of V.U.N. and ratio performance. A study of the effect of these parameters on sampling performance can be found in H.5.

### v2

@@ 5.1 Generative Performance Evaluation
Table 1 reports the results on two synthetic datasets: ER-DAG and SBM. In both settings, Directo consistently achieves the best trade-off between sample quality and validity. For ER-DAG, it achieves the highest validity score using directed RRWP as positional encoding, with a V.U.N. ratio of 94%. Notably, while LayerDAG enforces acyclicity by design, it fails to capture the ER structure of the target graph distribution, as evidenced by its low V.U.N. score (21.5%; see Table 16 for details). In the SBM case, Directo also successfully captures the distribution and consistently outperforms the baselines, achieving up to 99.5% in V.U.N. ratio and keeping low average graph statistics ratios. We provide the complete results for the distributional alignment metrics and the results for the other two synthetic datasets in Appendices H.1 and H.4.
@@
Table 1 shows the performance for the two real-world datasets. In TPU tiles, all Directo variants significantly outperform DAG generation baselines in terms of structure-aware metrics. The ratio scores for Directo are consistently and substantially lower than those observed in MLE, and especially LayerDAG, which indicates a more realistic distribution. Again, Directo variants achieve higher V.U.N. than all baselines, with the exception of LayerDAG, which benefits from enforcing acyclicity, the only validity constraint evaluated in this case. For Visual Genome, results reinforce the versatility of Directo. Directo RRWP achieves the highest V.U.N., indicating a superior ability to generate diverse and novel graphs while maintaining structural fidelity. Finally, we also see that Directo achieves the lowest RBF MMD score for both datasets, demonstrating that the model is successful in capturing the joint node-label dependencies of the distribution in the original dataset. Complete results are available in Tables 10 and 11 in Appendix H.2.
Conditional generation
As groundwork for directed graph generation, we prioritize strong unconditional generation, which we consider a key prerequisite for effective digraph modeling. At the same time, steering generation with graph-level properties is important for many downstream tasks. To illustrate the flexibility of our framework, we show that Directo can be extended to conditional generation without major architectural changes via classifier-free guidance (Ho & Salimans, 2022; Nisonoff et al., 2025). On the TPU Tiles dataset, we condition on the execution time of the computational graph, aiming at low execution times. Directo performs competitively with specialized autoregressive models like GraphRNN and LayerDAG (Appendix H.7), even surpassing the OneShotDAG variant of Li et al. (2025).
5.2 Ablation study
Impact of dual attention
@@ 5.2 Ablation study
We now analyze the influence of the dual attention mechanism on the performance of Directo. In Figure 3, we compare the V.U.N metric and the average graph statistics ratios (“Ratio”) score across both synthetic datasets (ER-DAG and SBM), highlighting the impact of dual attention on generation quality and graph realism. We observe that the dual attention mechanism consistently improves generative performance, regardless of the positional encoding it is combined with. Notably, even in the absence of any positional encoding (“No PE”), dual attention still achieves non-zero V.U.N., highlighting its capacity to independently capture directionality-relevant information. See Appendix H.3 for the full ablation tables, both for Directo and Directo-DD.
Impact of positional encodings
We also evaluate the sensitivity of Directo to different positional encodings, including a direction-agnostic baseline with the Laplacian (“Lap”) on synthetic datasets (see Appendix B for details on how it is computed using the symmetrized adjacency matrix). In Figure 3, we observe consistent trends across datasets: integrating positional encodings improves V.U.N. and Ratio, mirroring the undirected setting; moreover, direction-aware encodings outperform agnostic ones, supporting our modeling hypothesis. Among the encodings tested, Directed RRWP achieves the best overall performance on ER-DAG, and MultiMagLapPE with $Q=10$ on SBM. Nevertheless, we remark that RRWP demonstrates clear superiority in both scalability and runtime efficiency (see Section G.3). Overall, while positional encodings influence performance, the dual attention mechanism remains the most critical component for capturing directional dependencies. Full tables for both Directo and Directo-DD can be found in Appendix H.4.
@@
Table 2: Directed graph generation performance for the dual attention mechanism versus doubling the depth of the standard network.
{adjustbox}
width= Dataset Model Ratio $\downarrow$ V.U.N. $\uparrow$ ER-DAG RRWP - Double 4.9 $\pm$ 0.2 72.0 $\pm$ 9.8 RRWP - Dual 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 MagLap - Double 4.3 $\pm$ 0.7 80.0 $\pm$ 6.3 MagLap - Dual 1.3 $\pm$ 0.2 91.0 $\pm$ 2.5 SBM RRWP - Double 27.1 $\pm$ 2.3 0.0 $\pm$ 0.0 RRWP - Dual 1.4 $\pm$ 0.2 87.0 $\pm$ 5.1 MagLap - Double 24.8 $\pm$ 4.6 6.0 $\pm$ 4.9 MagLap - Dual 1.9 $\pm$ 0.3 77.0 $\pm$ 7.6
Scalability study
We investigate scalability across three axes: dataset size, parameter efficiency, and graph size (Appendix H.6). Increasing the amount of training data improves performance but comes with longer training time (Table 17). For parameter efficiency, our dual attention mechanism clearly demonstrates better performance than simply enlarging the adapted undirected model, as it can be seen in Table 2. Finally, while larger graphs pose challenges, particularly for enforcing strict acyclicity, Directo achieves strong generation quality (Ratio) across all graph sizes commonly used in the literature for the applications explored in this paper (Table 20).
@@ 6 Related work
Initial methods include auto-regressive models (You et al., 2018; Liao et al., 2019), which grow the graph progressively by inserting nodes and edges, but require node ordering for feasibility. On the other hand, one-shot graph generative models such as VAEs (Kipf & Welling, 2016; Simonovsky & Komodakis, 2018; Jin et al., 2018; Ma & Zhang, 2021; Vignac & Frossard, 2022), GANs (De Cao & Kipf, 2018; Krawczuk et al., 2021; Martinkus et al., 2022), and normalizing flows (Kaushalya et al., 2019; Wehenkel & Louppe, 2021; Lippe & Gavves, 2021; Luo et al., 2021b) can directly generate full graphs in a single forward pass, thereby eliminating the need for predefined node ordering. Graph diffusion models also belong to this class, with the initial approaches consisting of adaptations of continuous state-space discrete-time diffusion frameworks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019) to the graph setting. These include works such as EDP-GNN (Niu et al., 2020), DGSM (Luo et al., 2021a), and GeoDiff (Xu et al., 2022). Later, the discrete state-space diffusion framework (Austin et al., 2021) was adopted, with DiGress (Vignac et al., 2023a) and MCD (Haefeli et al., 2022) operating in discrete time. On the other hand, methods that operate in continuous time (Song et al., 2021; Campbell et al., 2022) were introduced, including GDSS (Jo et al., 2022) and GruM (Jo et al., 2024) in continuous state-spaces and DisCo (Xu et al., 2024) and Cometh (Siraudin et al., 2024) for discrete state-spaces. Recent approaches combine the sequential modeling of auto-regressive methods with the global denoising of diffusion (Zhao et al., 2024). Finally, DeFoG (Qin et al., 2025a) recently achieves state-of-the-art performance by leveraging discrete flow matching (Campbell et al., 2024; Gat et al., 2024).
@@
The vast majority of works in this area focus solely on DAG generation, with research including autoencoders (Zhang et al., 2019), autoregressive methods (Li et al., 2022) and, more recently, combining diffusion in different steps of an auto-regressive process (Li et al., 2025). However, they require the input DAGs to be topologically ordered, which can result in higher computational cost. Other approaches (Asthana et al., 2024; An et al., 2024) use ideas from discrete diffusion for Neural Architecture Search (NAS), but they are tailored to the specific application and, therefore, to DAGs. Finally, Law et al. (2025) are the first to propose a general digraph generation method from an auto-regressive perspective. Unfortunately, to the best of our knowledge, its implementation is not publicly available, limiting its accessibility for study and evaluation.
@@ 7 Conclusion and future directions
We propose Directo, a discrete flow matching-based method for directed graph generation. It combines directionality-aware positional encodings with a dual-attention mechanism that handles both source-to-target and target-to-source dependencies, overcoming the limitations of naïve extensions from undirected models. We also introduce a dedicated benchmarking framework to address the lack of standardized evaluation. Empirical results show that Directo outperforms existing baselines on synthetic and real-world datasets, successfully generating directed structures and preserving key properties, such as acyclicity, without explicit constraints, highlighting its generality and robustness.
Limitations and future directions
Despite these promising results, limitations remain. First, scalability can be further improved in handling larger datasets and graph sizes (a direct result of the combinatorial complexity of digraphs). Incorporating scalable strategies, such as sparsity-aware attention (Qin et al., 2025b), hierarchical (Bergmeister et al., 2024; Jang et al., 2024), or latent diffusion (Yang et al., 2024), could help extend Directo at scale. Second, although our current setup includes classifier-free conditional generation, exploring alternative conditional strategies could further improve performance. Third, while Directo can learn structural properties like acyclicity implicitly, enforcing such constraints (e.g., via PRODIGY (Sharma et al., 2024) or ConStruct (Madeira et al., 2024)) could further improve controllability and domain-specific validity. Exploring these directions would enhance the applicability and impact of digraph generative models, particularly on new real-world scenarios. Appendix K includes a detailed discussion of limitations and future work.
@@ Appendix A Dual Attention
To incorporate edge features into the attention mechanism, we apply FiLM-style modulation (Perez et al., 2018) using additive and multiplicative projections $\bm{W}_{\text{mul}},\bm{W}_{\text{add}}$:
@@
where $X_{\to y}$ and $E_{\to y}$ are PNA layers (Corso et al., 2020) that given node features $\bm{X}\in\mathbb{R}^{B\times N\times d_{x}}$ or edge features $\bm{E}\in\mathbb{R}^{B\times N\times N\times d_{e}}$ and a learnable weight matrix $\bm{W}_{\text{PNA}}$, computes
@@ Appendix B Positional Encodings
To retain directional information while preserving desirable spectral properties, Geisler et al. (2023) propose the Magnetic Laplacian, which introduces a complex-valued phase encoding into the adjacency matrix. Specifically, the (unnormalized and normalized) Magnetic Laplacians are given by:
@@
Recently,  Huang et al. (2025) extends the Magnetic Laplacian by introducing a multi-$q$ formulation, which considers as positional encoding stacking together the eigenvalues and eigenvectors of $Q$ distinct Magnetic Laplacians with potentials $q_{1},\ldots,q_{Q}$. This approach enables the recovery of the bidirectional walk profile, a generalization of walk counting in undirected graphs that captures a broader range of directional relationships.
@@
For directed graphs, it is possible to consider $\bm{T}=\bm{A}\bm{D}^{-1}_{out}$, where we take into account the degree matrix of the outgoing edges $\bm{D}^{-1}_{out}$ (Geisler et al., 2023). The $K$-step transition probabilities are then captured by powers of $\bm{T}$, producing the sequence $\bm{I},\bm{T},\bm{T}^{2},\ldots,\bm{T}^{K-1}]$.
Additionally, in the directed setting,  Geisler et al. (2023) suggest it is also informative to model reverse random walks, starting from incoming edges. For this, we can define the reverse transition matrix $\bm{R}=\bm{A}^{\top}\bm{D}_{\text{in}}^{-1}$, where $\bm{D}_{\text{in}}$ is the in-degree matrix. This would represent the likelihood of arriving at a given edge starting from another through $k$-step random walks. Analogously, we compute $\bm{I},\bm{R},\bm{R}^{2},\ldots,\bm{R}^{K-1}]$.
To have our final positional encoding, following Geisler et al. (2023), we ensure that graphs are not nilpotent and that the probabilities encoded in the matrices sum up to 1 by adding self-loops to sink nodes during pre-processing. Finally, the full positional encoding concatenates both forward and reverse walk features for a pre-defined $K$, yielding:
@@ Appendix C Sampling Optimization in Directo
The decoupling of training and sampling in discrete flow matching enables principled, training-free modifications to the sampling procedure, offering a flexible design space to enhance generation quality across digraph distributions (Qin et al., 2025a). In this section, we detail the three core parameters that govern sampling behavior: (i) time distortion functions, (ii) target guidance intensity, and (iii) stochasticity strength. Each plays a distinct role in shaping the denoising trajectory and contributes to different aspects of generation quality and structural fidelity.
@@
We consider the same five representative time distortion functions from Qin et al. (2025a) that yield diverse distributions for $t^{\prime}$:
@@
The final sampling-time parameter is the stochasticity coefficient $\eta$. This parameter scales an auxiliary rate matrix $R_{t}^{\text{DB}}$ that satisfies the detailed balance condition (Campbell et al., 2024) and adds it to the optimal rate matrix $R_{t}^{*}$:
@@
Together, these sampling parameters constitute a flexible, principled toolkit for tailoring the sampling procedure to diverse graph domains and structural requirements. In our case, we optimize these parameters at sampling time to improve the performance of Directo. In particular, we individually optimize each of them for the main results, and perform the final sampling using the best performing parameters in terms of V.U.N (a detailed description is available in Appendix G.2). and ratio performance. A study of the effect of these parameters on sampling performance can be found in H.8.